OpenStudy (anonymous):

given functions...pls. compute for: a) (fog)(x) b) (fog)(-2) 1.) f(x)=x2-2;g(x)=x3

OpenStudy (anonymous):

\[f\circ g(x)=f(g(x))=f(x^3)=(x^3)^2-2=x^6-2\]

OpenStudy (anonymous):

so \[f\circ g(-2)=(-2)^6-2=64-2=62\]